# SurgWorld: Learning Surgical Robot Policies from Videos via World Modeling

### v2

@@ 1 Introduction
In parallel, the robotics community has seen tremendous progress in large vision–language–action (VLA) models. Models such as RT-2 [brohan2023rt2], OpenVLA [liu2024openvla], and GR00T [bjorck2025gr00t] has demonstrated the potential of foundation models to generalize across diverse robotic manipulation tasks. These models leverage large-scale multimodal datasets that couple visual observations, language descriptions, and robot actions, enabling rich world understanding and robust policy learning, yielding significantly improved generalization to novel objects and commands in non-surgical manipulation domains. Similarly, recent work on synthetic data generation such as DreamGen [jang2025dreamgen] shows that video world-models combined with inverse dynamics can produce synthetic paired video–action datasets that boost policy learning in general manipulation tasks.
@@ 2 Related Work
VLA Models and Imitation Learning (IL). Large-scale vision–language–action (VLA) models have recently emerged as a powerful paradigm for general-purpose robotic policy learning. Recent open-source efforts such as OpenVLA [liu2024openvla], $\pi_{0}$ [black2410pi0], and GR00T N1 [bjorck2025gr00t] have been trained on massive and diverse data sources, leveraging diverse robot embodiments and multi-task datasets and can robustly handle real-world variability. To train those VLAs, imitation learning (IL) remains the most direct and scalable approach, typically via behavior cloning (BC) [pomerleau1989alvinn]. However, BC suffers from covariate shift, producing rapidly worsening performance when the dataset is limited [ross2011dagger]. In robotics, several works aim to improve data efficiency using offline datasets [bahdanau2019offline, kumar2020offline] or synthetic augmentation [james2019sim2real]. However, despite these advances, current VLA models still rely heavily on imitation learning on large scale paired image–action datasets, which are scarce in specialized domains such as surgical robotics.
@@
Learning Policy Models from Videos. Learning policy models directly from video has become increasingly important due to the lack of datasets with kinematics. Ye et allet@tokeneonedot [ye2024latent] and Jang et allet@tokeneonedot [jang2025dreamgen] propose a latent-action pretraining scheme using internet-scale videos without robot action labels to bootstrap vision-language-action models. Hu et allet@tokeneonedot [hu2025videoprediction] train predictive visual representations via video diffusion models and embed an implicit inverse-dynamics policy conditioned on these representations. Bharadhwaj et allet@tokeneonedot [bharadhwaj2025gen2act] use human-video generation to condition robot policies on novel scenarios, reducing reliance on robot-collected data. Li et allet@tokeneonedot [li2025unifiedvideoaction] present a unified video-action latent model that jointly handles forward/inverse dynamics, video generation, and action inference within one framework. Tian et allet@tokeneonedot [tian2025predictiveinverse] close the loop by using inverse-dynamics models conditioned on predicted visual states for large-scale manipulation training. These developments directly inspire our approach, which uses IDM to connect world models with surgical robot learning.
Automated Surgical Robotics. Automation in surgical robotics have increasingly leveraged learning-based approaches. Long et allet@tokeneonedot [long2025surgical] introduced a vision-based embodied intelligence framework that enables zero-shot sim-to-real transfer for a variety of laparoscopic assistive tasks using imitation learning, Kim et allet@tokeneonedot [kim2024surgical] proposed the Surgical Robot Transformer (SRT), which addresses the da Vinci robot’s inaccurate kinematics by formulating actions in a hybrid-relative space, achieving high success rates on fundamental tasks. The same group later developed SRT-H [kim2025srt] that enables step-level autonomy in complex procedures such as cholecystectomy. While these works demonstrate impressive progress in task autonomy, they also reveal a critical need for large-scale, diverse surgical demonstration datasets, which in surgery is uniquely challenging: obtaining large-scale paired visual–kinematic datasets requires specialized hardware and surgeon supervision.

### v3

@@ 1 Introduction
In parallel, the robotics community has seen tremendous progress in large vision–language–action (VLA) models. Models such as RT-2 [brohan2023rt2], OpenVLA [kim2024openvla], and GR00T [bjorck2025gr00t] has demonstrated the potential of foundation models to generalize across diverse robotic manipulation tasks. These models leverage large-scale multimodal datasets that couple visual observations, language descriptions, and robot actions, enabling rich world understanding and robust policy learning, yielding significantly improved generalization to novel objects and commands in non-surgical manipulation domains. Similarly, recent work on synthetic data generation such as DreamGen [jang2025dreamgen] shows that video world-models combined with inverse dynamics can produce synthetic paired video–action datasets that boost policy learning in general manipulation tasks.
@@ 2 Related Work
VLA Models and Imitation Learning (IL). Large-scale vision–language–action (VLA) models have recently emerged as a powerful paradigm for general-purpose robotic policy learning. Recent open-source efforts such as OpenVLA [kim2024openvla], $\pi_{0}$ [black2410pi0], and GR00T N1 [bjorck2025gr00t] have been trained on massive and diverse data sources, leveraging diverse robot embodiments and multi-task datasets and can robustly handle real-world variability. To train those VLAs, imitation learning (IL) remains the most direct and scalable approach, typically via behavior cloning (BC) [pomerleau1989alvinn]. However, BC suffers from covariate shift, producing rapidly worsening performance when the dataset is limited [ross2011dagger]. In robotics, several works aim to improve data efficiency using offline datasets [kumar2020offline] or synthetic augmentation [james2019sim2real]. However, despite these advances, current VLA models still rely heavily on imitation learning on large scale paired image–action datasets, which are scarce in specialized domains such as surgical robotics.
@@
Learning Policy Models from Videos. Learning policy models directly from video has become increasingly important due to the lack of datasets with kinematics. Ye [ye2024latent] and Jang [jang2025dreamgen] propose a latent-action pretraining scheme using internet-scale videos without robot action labels to bootstrap vision-language-action models. Hu [hu2025videoprediction] train predictive visual representations via video diffusion models and embed an implicit inverse-dynamics policy conditioned on these representations. Bharadhwaj [bharadhwaj2025gen2act] use human-video generation to condition robot policies on novel scenarios, reducing reliance on robot-collected data. Li [li2025unifiedvideoaction] present a unified video-action latent model that jointly handles forward/inverse dynamics, video generation, and action inference within one framework. Tian [tian2025predictiveinverse] close the loop by using inverse-dynamics models conditioned on predicted visual states for large-scale manipulation training. These developments directly inspire our approach, which uses IDM to connect world models with surgical robot learning.
Automated Surgical Robotics. Automation in surgical robotics have increasingly leveraged learning-based approaches. Long [long2025surgical] introduced a vision-based embodied intelligence framework that enables zero-shot sim-to-real transfer for a variety of laparoscopic assistive tasks using imitation learning, Kim [kim2024surgical] proposed the Surgical Robot Transformer (SRT), which addresses the da Vinci robot’s inaccurate kinematics by formulating actions in a hybrid-relative space, achieving high success rates on fundamental tasks. The same group later developed SRT-H [kim2025srt] that enables step-level autonomy in complex procedures such as cholecystectomy. While these works demonstrate impressive progress in task autonomy, they also reveal a critical need for large-scale, diverse surgical demonstration datasets, which in surgery is uniquely challenging: obtaining large-scale paired visual–kinematic datasets requires specialized hardware and surgeon supervision.